OpenStudy (hang254):

Express the complex number in trigonometric form. -2 + 2sqrt3 i

OpenStudy (hang254):

\[-2+2\sqrt{3}i\]

OpenStudy (hang254):

@jim_thompson5910 @aaronq

OpenStudy (anonymous):

you need two numbers \(r\) and \(\theta\) in order to write this as \[-2+3\sqrt 3 i=r\left(\cos(\theta)+i\sin(\theta)\right)\]

OpenStudy (anonymous):

\(r\) is the easiest to find, is is \[r=\sqrt{a^2+b^2}\] in your case \(a=-2,b=2\sqrt3\)

OpenStudy (hang254):

ok, lets see

OpenStudy (anonymous):

k i will for you let me know what you get

OpenStudy (hang254):

1.414

OpenStudy (anonymous):

forget the calculator

OpenStudy (anonymous):

just do it by hand

OpenStudy (hang254):

4

OpenStudy (anonymous):

\[\sqrt{2^2+(2\sqrt3)^2}=\sqrt{4+12}=\sqrt{16}=4\] right

OpenStudy (hang254):

alright

OpenStudy (anonymous):

now you need \(\theta\) which may or may not be obvious

OpenStudy (anonymous):



OpenStudy (anonymous):

you can use \(\tan(\theta)=\frac{b}{a}\)

OpenStudy (anonymous):

or in your example \[\tan(\theta)=-\sqrt3\]

OpenStudy (hang254):

ok, i'm sorry for wasting your time, i actually solved this problem already and mistakenly didn't mark it off as complete

OpenStudy (hang254):

the answer is 4 (cos(2pi/3) + i sin (2pi/3) sorry again, thanks!

OpenStudy (hang254):

@satellite73

OpenStudy (anonymous):

lol yw